OpenStudy (anonymous):

Hi. I am new. I need help with this math question. I just need explaining and guidance on how to do it. It is a solving system question 2x = 5y + 4 3x - 2y = -16 @PotatoOfHistory @SolomonZelman

OpenStudy (anonymous):

You can use the substitution method but do they tell you what method they want you to use?

OpenStudy (anonymous):

No that is the reason why I am having a hard time

OpenStudy (anonymous):

Do you know how to do this problem?

OpenStudy (anonymous):

@MDoodler

OpenStudy (anonymous):

Here is one way. Take 2x = 5y + 4 and put it into standard form, which would be like 2x = 5y + 4 2x -5y = 4 Now we have 2x -5y = 4 3x - 2y = -16 Can you finish this yet?

OpenStudy (anonymous):

Here's what I got for the next step. 3(2x-5y) - 2y = -16 Then I distribute the 3x right?

OpenStudy (anonymous):

THANK YOU SO MUCH! I really appreciate the in depth answer. It was my mistake to not give you the answer choices to help me. I am sorry. I am new at this. But your answer doesn't match the answer choices I was given. Here they are: A: (-8, -4) B: 8,4 C: (-4,-8) D: (4,8)